英語 (English)

発信型英語 (Communicative English) (工 ((建 A ー 1)2 年))

座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部 2 単位後期火 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 This Communicative English class planned to give a good assistance

in rising up English speaking and communicating abilities. 【授業の概要】 The Classes will be running in ENGLISH language only for more

practices chance. There is no specific text book but there will be some other media depending on the ’Topics’ which will be selected up to date. Speaking out practices will take place during the classes.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】To create a person with so high self-confidence and effective

communication skills. 【授業の計画】

1. Overall Guidance

2. Conversational Topic 1

3. Conversational Topic 2

4. Conversational Topic 3

5. Conversational Topic 4

6. Conversational Topic 5

7. Conversational Topic 6

8. Conversational Topic 7

9. Conversational Topic 8

10. Conversational Topic 9

11. Conversational Topic 10

12. Conversational Topic 11

13. Conversational Topic 12

14. Conversational Topic 13

15. Conversational Topic 14

16. Review 【教科書】There is no specific text book but there will be some other media

depending on the ’Topics’ which will be selected up to date.

【参考書等】 [参考資料]

【成績評価の方法】 Attendance, Homework fulfillment and Participation in class 【再試験の有無】 There is no make-up examination for this course. 【受講者へのメッセージ】 Asking questions in class is very important.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221214 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ 座喜 .

English

Communicative English (工 ((建 A ー 1)2 年))

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences 2 units 後期火 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This Communicative English class planned to give a good assistance in

rising up English speaking and communicating abilities. Outline⟩ The Classes will be running in ENGLISH language only for more

practices chance. There is no specific text book but there will be some other media depending on the ’Topics’ which will be selected up to date. Speaking out practices will take place during the classes.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ To create a person with so high self-confidence and effective communication

skills.

Schedule⟩

1. Overall Guidance

2. Conversational Topic 1

3. Conversational Topic 2

4. Conversational Topic 3

5. Conversational Topic 4

6. Conversational Topic 5

7. Conversational Topic 6

8. Conversational Topic 7

9. Conversational Topic 8

10. Conversational Topic 9

11. Conversational Topic 10

12. Conversational Topic 11

13. Conversational Topic 12

14. Conversational Topic 13

15. Conversational Topic 14

16. Review Textbook⟩ There is no specific text book but there will be some other media

depending on the ’Topics’ which will be selected up to date.

Reference⟩ [参考資料]

Evaluation Criteria⟩ Attendance, Homework fulfillment and Participation in class Re-evaluation⟩ There is no make-up examination for this course. Message⟩ Asking questions in class is very important.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221214 Contact (Office-Hour, Room, E-mail)⟩

⇒ Zaki .